Work out the following and give answers in the indicated bases/codes. (Show your ALL working):

a. 16608 – 53016 in base 8

b. AB.7516 – 830.510 in base 2

c. 402810 + 475.0816 in base 16

d. 102410 – 110101112 in base 2 


1
Expert's answer
2022-05-04T17:42:32-0400

a) "1660_8-530_{16}=1660_8-2460_8=-600_8"

"530_{16}=(0101)_2(0011)_2(0000)_2=10100110000_2=(10)(100)(110)(000)_2=2460_8"


b) "\text{AB}.75_{16}-830.5_{10}=1010100000110110.1_2"

"\text{AB}.75=\text{10}_{10}\text{11}_{10}.7_{10}5_{10}=1010_{2}1011_20111_20101_2=1010101101110101_2"

"830.5_{10}=2\cdot 415+2^{-1}=2\cdot207+2+2^{-1}=2^3\cdot 103+2^2+2+2^{-1}=2^4\cdot 51+2^3+2^2+2+2^{-1}=2^5\cdot 25+2^4+2^3+2^2+2+2^{-1}=2^8\cdot 3+2^5+ 2^4+2^3+2^2+2+2^{-1}=2^9+2^8+2^5+
2^4+2^3+2^2+2+2^{-1}=1100111110.1_2"


c) "4028_{10}-475.08_{16}=\text{B}46.\text{F}8_{16}"

"4028_{10}=16\cdot 251+12=16\cdot(16\cdot 15+11)+12=16^2\cdot 15+16\cdot 11+12=\text{FBC}_{16}"


d) "1024_{10}-11010111_2= 10000000000_2-11010111_2=1100101001_2"

"1024_{10}=2^{10}=10000000000_2"